OpenStudy (anonymous):

how to work out 2 x 1/3

OpenStudy (usman1995):

apply DMAS rule

OpenStudy (usman1995):

division,multiplication, addition subtraction follow the order

OpenStudy (usman1995):

it's simple 2/3

OpenStudy (anonymous):

Just start with 2 x 1 then do the / 3 2 x 1 / 3 = 2 / 3 = Then put the answer as either \[\frac{ 2 }{ 3 }\] or 0.666)

OpenStudy (anonymous):

its quite simple.. you can just multiply 2 with 1 that is 2 and then divide it with 3. so your answer becomes 2/3. :)